OpenStudy (anonymous):

given the proportion a/b = 8/15, what ratio completes the equivalent proportion a/8 = ? the answer is b/15 , but how do you get it?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ a }{ b }=\frac{ 8 }{ 15 }\]First multiply b on both sides\[\frac{ a }{ b }~\times~b=\frac{ 8 }{ 15 }~\times~b\]\[a=\frac{ 8b }{ 15 }\]Then divide 8 on both sides \[a~\div~8=\frac{ 8b }{ 15 }~\div~8\]\[\frac{ a }{ 8 }=\frac{ b }{ 15 }\]

OpenStudy (anonymous):

@MARC_ Thank you so much!

OpenStudy (anonymous):

yw!